OpenStudy (anonymous):

How would I write this in equation form?? Tom has a collection of 30 CDs and Nita has a collection of 18 CDs. Tom is adding 1 CD per month to his collection while Nita is adding 5 CDs per month to her collection. After how many months will Tom and Nita have the same number of CDs?

OpenStudy (anonymous):

equation to show work: 1x + 30 = ? 5x + 18 = ? x = the number of moths x = 3

OpenStudy (texaschic101):

m stands for # of months 30 + 1m = 18 + 5m now solve for m to see how many months they will be the same

OpenStudy (anonymous):

and do I combine like terms? or where do I start

OpenStudy (anonymous):

@texaschic101